OpenStudy (anonymous):

prove cot-1 x = pi/2 - tan -1 x

OpenStudy (anonymous):

could u plz draw out the equation?

OpenStudy (solomonzelman):

\[\cot^{-1}x=\frac{π}{2}-\tan^{-1}x~~~~~~~~~~~~~~~~~~this?\]

OpenStudy (anonymous):

yes as solomon put it

OpenStudy (anonymous):

@SolomonZelman my brain won't function can u do this one xD

OpenStudy (solomonzelman):

\[π=180°\]

OpenStudy (anonymous):

still waiting for help :)

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{1}{Tan^{-1}x} +Tan^{-1}x } =90\] \[\huge\color{red}{ Let~~~~Tan^{-1}x=a } \] \[\huge\color{red}{ \frac{1}{a}+a=90 } \]\[\huge\color{red}{ 1+a^2=90a } \]\[a^2-90a+1=0\]

OpenStudy (solomonzelman):

\[\huge\color{red}{ \frac{90±\sqrt{8100-4} }{2} } \] \[\huge\color{red}{ \frac{90±94.95}{2} } \] \[\huge\color{red}{ 45±47.475 } \]

OpenStudy (solomonzelman):

\[\huge\color{red}{ Tan^{-1}x=-2.457 } \]\[\huge\color{red}{ Tan^{-1}x = +92.475 } \]

OpenStudy (solomonzelman):

But this thing is so stupid though....

OpenStudy (raden):

cot-1 x = pi/2 - tan -1 x put cot on both sides, get cot(cot-1 x) = cot(pi/2 - tan -1 x) x = tan(tan-1 x) x = x :)

OpenStudy (anonymous):

thanks solomon and Rad!! :D